zeinananan:

Draw the graph of the function f(x)=int(x)+1 for the domain –2 ≤ x ≤ 2. (step functions). Someone please help me with this. I really need to submit this assignment right now. @vocaloid @dude

Vocaloid:

 this is int(x) (notice how each output value is rounded down to the nearest integer, so x = 1.5 gives y = 1 as the output) your function is int(x) + 1 so you just need to shift the graph up 1 unit

Vocaloid:

make sure to stay within the domain –2 ≤ x ≤ 2

zeinananan:

But i dont know where to stop. like where the domain ends. The end point would be on (2,3)?

Vocaloid:

yes

zeinananan:

You sure? Also, thank you so much !

Vocaloid:

yes it would start on (-2,-1), and go up to (2,3) following the same pattern in the graph, just be careful about closed/open circles and you should be good to go

zeinananan:

The left point should be closed and the right open. right?

Vocaloid:

yes

zeinananan:

The (2,3) point should be filled in or not?

Vocaloid:

yes

zeinananan:

hey, can u help me # solve one more question? its: A carpenter can build a toolshed three times as fast as his apprentice. Working together, they can build the shed in 6 hours. How long would it take each of them working alone to build the shed?

Vocaloid:

I've never been good at these but here's my best attempt if it takes them 6 hours together to complete the whole project, then in one hour, together, they do 1/6 of the total job we can represent the work done by the carpenter in 1 hour as x, and the work done by the apprentice in 1 hour as 3x so x + 3x = 1/6 solving for x gives x = 1/24. meaning the carpenter does 1/24 of the whole job per hour, meaning the whole job takes him 24 hours, alone the apprentice works 3 times as fast so the apprentice gets it done in 8 hours, alone and that's it

zeinananan:

YES THANKYOU